whitemonsterbunny17 (whitemonsterbunny17):

Help with a Geometry problem, please? Thanks! ^.^

whitemonsterbunny17 (whitemonsterbunny17):

If AB= 5x-8 and EF= 3x, then x= _____ ?

whitemonsterbunny17 (whitemonsterbunny17):

How would I set up the equation?

OpenStudy (1995monster):

Does AB=EF?

whitemonsterbunny17 (whitemonsterbunny17):

Yes.

OpenStudy (1995monster):

so 5x-8=3x

whitemonsterbunny17 (whitemonsterbunny17):

Okay, so then would I add the # of x's to get 8x?

OpenStudy (1995monster):

2x-8=AB you cant really solve for X because there are two X's in the equations

OpenStudy (1995monster):

Idk this is weird

jimthompson5910 (jim_thompson5910):

Whitemonsterbunny17, you can only say 5x+3x = 8x only if the x terms are on the same side

whitemonsterbunny17 (whitemonsterbunny17):

Huh? I have to solve for x though?

jimthompson5910 (jim_thompson5910):

but the x terms are on different sides

whitemonsterbunny17 (whitemonsterbunny17):

Oh, then what would I do? o.O

jimthompson5910 (jim_thompson5910):

you have to move the x terms so they all end up on the same side

jimthompson5910 (jim_thompson5910):

so either move the 5x over to the right side or move the 3x to the left side

whitemonsterbunny17 (whitemonsterbunny17):

Does it matter which side?

jimthompson5910 (jim_thompson5910):

nope

whitemonsterbunny17 (whitemonsterbunny17):

Okay, so then I add the x's to get 8x, right?

jimthompson5910 (jim_thompson5910):